OpenStudy (anonymous):

Find f. f ''(x) = 5 + cos x, f(0) = −1, f(7π/2) = 0

OpenStudy (anonymous):

f'(x)=5x+sin x + C f(x)=?x^2-cos x+Cx+D

OpenStudy (blurbendy):

hope this helps

OpenStudy (loser66):

f" (x) = 5 +cos x \[f'(x)= \int (5 + cosx)dx\]yea, he is right

OpenStudy (anonymous):

That is the answer. Thanks so much. It was very helpful to see the work

OpenStudy (blurbendy):

yw